The physician orders foscarnet 40 mg q.8h. IV push for a patient who weighs 50 kg. The drug reference manual states that the direct IV rate of administration for foscarnet is not to exceed 1 mg / kg over 1 min. The nurse administers the ordered dose over 48 seconds. Evaluate the nurse’s decision about the rate of administration.

Choose one below:

The nurse administered the dose of foscarnet too slowly.
The nurse has correctly calculated the rate of administration.

The nurse should have administered the dose over 1 minute.

The nurse should have administered the dose over 50 seconds.
Answer #1

ANSWER.

* Physicians order = 40 mg IV q8h.

    Patient weight = 50 kg.

   Reference range = 1 mg/kg over 1 minute.

    So

       Total amount (mg) can given to the patient over

                                                 one minute = 50 mg over 1 minute.

      Total amount nurse can administer over 48 sec is

                                                           1 min (60 sec) = 50 mg.

                                                                        1 sec = X mg.

                                                                         60X = 50 * 1

                                                                             X = 50 * 1 / 60

                                                                      1 sec = 0.83 mg.

              Amount nurse can administer for 48 sec = 0.83 * 48

                                                            = 39.98 mg.

                                                                                = 40 mg.

      * The nurse has correctly calculated the rate of administration.